Abstract for The field generated by the points of small prime order on an elliptic curve

We prove that the field generated by the points of order p of an elliptic curve over Q̄ is strictly bigger than the cyclotomic field generated by pth roots of unity, for p a prime number satisfying p = 11 or 17 ≤ p < 1000.